Simplify the following boolean algebra equation only with the xnor logic gate!

Please answer clearly

a. ABCD+ABCD+ACBD+ADBC+BCAD+BDAC +CDAB+ABCD b. B(A+CA)+A’(B’+(BD)+(BD)A)
Answer #1

Giren I am considering A as MSB and D as LSB, L and the erder as ABCD. a) Simplification ( Implementation using XNOR gales. A

→ (COD) (408) + (COD) (105) It û ù the form of Xoy, given that X = COD Y = AOB X Oy = x+ + Kỹ = (COD) (106) + ( COD) (AOB] {X

6) B(A + CA) + Al(8+(8D) + (BD)) B(All+c)) + ! [ B +(80) (1 + n]] = AB + Al( B+(BD) - AB + AB+ B+D! = AB + All B +01